OpenStudy (anonymous):

Hey, I'm having issues with Hyperbola's.... I need to find the center of (x-2)^2/25 - (y-3)^2/36 = 1 Can anyone help?

OpenStudy (dumbcow):

center is (2,3)

OpenStudy (anonymous):

.... I kinda wanted someone to explain how they got the answer (:

OpenStudy (dumbcow):

the general equation for hyperbola: \[\frac{(x-h)^{2}}{a^{2}} - \frac{(y-k)^{2}}{b^{2}} =1\] center is (h,k)

OpenStudy (anonymous):

Tyvm (:

OpenStudy (anonymous):

I have one more question for you Mr. Cow.... you said that the formula is (x−h)^2/a2−(y−k)^2/b2=1 does this mean that if the formula is (x-2)^2/25 that h = 2?

OpenStudy (dumbcow):

yes correct http://tutorial.math.lamar.edu/Classes/Alg/Hyperbolas.aspx here is a nice resource to help you out

OpenStudy (anonymous):

Cool thank you very much, could you answer my Asymptotes question as well? You are very helpful (: